the ratio of incomes of two persons is 9:7 The difference is $200. What are their weekly incomes? I don't know how to set it up in a system of equations.

Answer by mananth(16946)

let them be Mr. A & Mr B
Mr. A wages = x
Mr. B wages =y
x/y = 9/7
x=(9y/7)
x-y =200
substitute x
(9y/7)-y=200
multiply equation by 7
9y-7y=7*200
2y=1400
/2
y=700
So x = 900
Their wages are 900 & 700